Write down value of acceleration for free falling body.
1
Expert's answer
2020-09-30T12:29:27-0400

Let us consider the forces acting on the falling body. If we neglect the force of air resistance, we consider only the gravitational force that is equal to

"F=G\cdot\dfrac{mM_{\oplus}}{(R_{\oplus} +h)^2} ,"where h is the height above the ground.

F=ma, so

"a=G\cdot\dfrac{M_{\oplus}}{(R_{\oplus} +h)^2}".

If h=0, the acceleration is "g=9.81\,\mathrm{m/s^2}."